# Automated event generation for S-wave quarkonium and leptonium production in NRQCD and NRQED

### v1

@@ 1 Introduction
The main motivation of this work is to develop, and ultimately provide the community with an all-encompassing tool for bound-state production studies in NRQCD and non-relativistic QED (NRQED) Caswell:1985ui . The goal is to deliver reliable, efficient, and fast automated state-of-the-art computations, with NLO accuracy based on an extended Frixione–Kunszt–Signer (FKS) infrared-divergence subtraction formalism Frixione:1995ms ; Frixione:1997np ; AH:2024ueu planned for near-term implementation. This work lays the foundation for that endeavour and falls within the aforementioned category of process-independent parton-level event generators, designed to simulate processes involving one or more quarkonia and/or leptonia. As mentioned, our implementation is embedded within the well-established MG5_aMC framework, which is widely used in the high-energy physics community, thereby extending its capabilities to include S-wave quarkonium and leptonium production. To the best of our knowledge, no other publicly available tool currently provides similar functionality for leptonium processes.
@@ 2 Theoretical framework
The short-distance cross section ${\rm d}\hat{\sigma}({\cal I}_{1}{\cal I}_{2}\rightarrow(C_{1}C_{2})[n]+X_{p})$ can be obtained from the production amplitude of $C_{1}$ and $C_{2}$ by projecting it onto the specific quantum state $n$, following section 2 of ref. AH:2024ueu . As mentioned previously, the quantum state $n$ is specified by its $S$, $L$, $J$ and $C$ quantum numbers. For leptonium ($\mathcal{B}=\mathcal{L}$), only colour singlet configuration ($C=1$) is possible.
@@ 5 Benchmark processes and validation
In this section, we benchmark our LO implementation by comparing its predictions for various observables across multiple processes in $e^{+}e^{-}$ and $pp$ collisions with the corresponding results from HELAC-Onia Shao:2012iz ; Shao:2015vga . The comparison encompasses processes with single and multiple quarkonium final states, as well as quarkonium production in association with SM elementary particles. Our aim is to validate our calculations and assess their consistency with other established tools. This benchmarking provides an important cross check, ensuring the reliability of our predictions for the applications discussed in the following two sections.
@@ 6 Quarkonium production
To highlight the broad applicability of our implementation across different collision systems, we present predictions for quarkonium production in proton-proton ($pp$) collisions (section 6.1), electron-proton ($e^{-}p$) collisions (section 6.2), and electron-positron ($e^{+}e^{-}$) collisions (section 6.3).
@@
We emphasise that the broad range of processes considered in what follows is intended to demonstrate the versatility of our implementation; the numerical predictions presented should not be directly compared with experimental measurements, where available. As previously stated, our current implementation is restricted to S-wave production at LO and does not, for instance, account for feeddown effects where they occur. Nevertheless, we provide a sufficiently detailed exposition of the results to allow the reader to reproduce all numerical values shown using MG5_aMC with the same parameter inputs. In addition, we make our analyses scripts available as auxiliary files in the supplemental material accompanying the arXiv submission.
@@ 6.1.4 $J/\psi$ production in association with an open charm-anticharm pair
The LHCb experiment has observed the associated production of $J/\psi+D$ LHCb:2012aiv and $\Upsilon+D$ LHCb:2015wvu , with results indicating that SPS contributions alone are insufficient to describe the data, thereby necessitating the inclusion of DPS effects. Motivated by this strong phenomenological relevance, numerous theoretical studies have been dedicated to these processes Berezhnoy:2015jga ; Likhoded:2015fdr ; Karpishkov:2019vyt ; Shao:2020kgj ; Chernyshev:2023qea .
@@
Figure 1: Transverse momentum ($p_{T,\mathcal{Q}}$) distributions for $J/\psi$ production in association with $c\bar{c}$ in $pp$ collisions at $\sqrt{s}=13$ TeV. Panels (b-d) correspond to different $J/\psi$ Fock states, while panel (a) shows the total contribution obtained by combining them. The blue curve represents the fixed-order LO (fLO) prediction from MG5_aMC, and the orange curve shows the MG5_aMC LO events combined with PS from Pythia v8.315 (LO+PS). Both include their respective 7-point scale-variation uncertainty bands. The lower panels display the ratio to the fLO contribution in each case.
@@
In figure 1, we compare the $J/\psi$ transverse momentum spectra from the fixed-order LO (fLO) calculation and the LO prediction matched to PS (LO+PS). The $p_{T,J/\psi}$ spectrum including all three Fock states is shown in figure 1(a), while figures 1(b)1(c), and 1(d) display the spectra for the individual $J/\psi\!\left[{}^{3}S_{1}^{[1]}\right$], $J/\psi\!\left[{}^{1}S_{0}^{[8]}\right$], and $J/\psi\!\left[{}^{3}S_{1}^{[8]}\right$] Fock states, respectively. We observe that PS effects broaden the $p_{T,J/\psi}$ distributions in the $J/\psi\!\left[{}^{3}S_{1}^{[1]}\right$] and $J/\psi\!\left[{}^{1}S_{0}^{[8]}\right$] channels, while they narrow the transverse-momentum distribution in the $J/\psi\!\left[{}^{3}S_{1}^{[8]}\right$] channel. These opposite shifts from the two spin-triplet Fock states partially compensate each other in the total (figure 1(a)), resulting in a moderate overall distortion of the $p_{T,J/\psi}$ spectrum. Unlike the processes $pp\to J/\psi\!\left[{}^{3}S_{1}^{[8]},\,^{1}S_{0}^{[8]},\,^{3}P_{J}^{[8]}\right]+j+X$, the cross sections for these processes at small $p_{T,J/\psi}$ are finite. In the former case, one would need to impose cuts, and the effect of showering in presence of such cuts would be more significant Lansberg:2024zap .
@@ 6.1.5 Inclusive quarkonium-pair production
In contrast, $J/\psi+\Upsilon$ production in hadron-hadron collisions has now been measured at both the Tevatron by D0 D0:2015dyx and at the LHC by LHCb LHCb:2023qgu . The main motivation for studying charmonium plus bottomonium production is to probe colour-octet quarkonium production mechanisms Ko:2010xy ; Lansberg:2015lva ; Lansberg:2020rft , DPS Likhoded:2015zna ; Shao:2016wor , and potential new exotic hadrons such as fully-heavy tetraquarks ($b\bar{b}c\bar{c}$). Unlike $J/\psi$-pair or $\Upsilon$-pair production, the SPS colour-singlet process $pp\to J/\psi\!\left[{}^{3}S_{1}^{[1]}\right]\Upsilon\!\left[{}^{3}S_{1}^{[1]}\right]+X$ vanishes in QCD at LO ($\mathcal{O}(\alpha_{s}^{4}$)) due to quantum-number conservation. In fact, the lowest-order contribution in QCD is loop-induced at $\mathcal{O}(\alpha_{s}^{6})$ Shao:2016wor . This suppression of the SPS colour-singlet channel makes DPS contributions more prominent and increases the relative importance of colour-octet mechanisms in the SPS channel.
@@
Table 9 shows our predictions for the production of charmonium or bottomonium pairs, while table 10 presents the production of mixed charmonium–bottomonium pairs as well as $B_{c}^{\pm}$ pairs. All cross sections are computed at $\mathcal{O}(\alpha_{s}^{4})$. For charmonium or bottomonium pairs, the cross sections are dominated by colour-singlet configurations when both quarkonia are of the same type, e.g., in the production of two $\eta_{c}$ ($\eta_{b}$) or two $J/\psi$ ($\Upsilon$) mesons. By contrast, at LO in QCD, processes involving one pseudoscalar and one vector quarkonium of the same flavour, such as $\eta_{c}+J/\psi$ ($\eta_{b}+\Upsilon$), are typically dominated by configurations involving at least one colour-octet state. This is because, due to C parity, the pure colour-singlet partonic channels $gg\to(Q\bar{Q})\!\left[{}^{1}S_{0}^{[1]}\right](Q\bar{Q})\!\left[{}^{3}S_{1}^{[1]}\right$] and $q\bar{q}\to(Q\bar{Q})\!\left[{}^{1}S_{0}^{[1]}\right](Q\bar{Q})\!\left[{}^{3}S_{1}^{[1]}\right$] are forbidden in QCD. This also explains why the cross sections for pseudoscalar-vector final states are smaller than those of the other two configurations in table 9.
@@
We again emphasise the broad physics potential of quarkonium-pair production studies for probing the interplay between SPS and DPS interactions. However, the predictions presented in this work account only for the SPS contribution at LO, using generic cuts and LDME values, for illustration purposes.
@@ 6.1.6 Inclusive triple $J/\psi$ production
To showcase the capabilities of our implementation, we consider triple $J/\psi$ production in $pp$ collisions. This serves as an illustrative example of a non-trivial three-quarkonium final state, proposed in refs. Shao:2019qob ; dEnterria:2016ids ; dEnterria:2017yhd as a probe of triple parton scattering (TPS), and recently measured by CMS CMS:2021qsn , marking the first experimental measurement of TPS and providing new insight into multi-parton scattering dynamics. The LO ($\mathcal{O}(\alpha_{s}^{6})$) total cross section in QCD at $\sqrt{s}=13\,\mathrm{TeV}$ is
@@ 6.2 Quarkonium production in electron-proton collisions
At electron-proton colliders, quarkonium production can occur via either DIS or photoproduction processes. In the following two subsections, we consider both types of collisions and compute cross sections for charmonium production. We retain the same SM input parameters as discussed in section 4, and here specify the relevant collider settings, which were not introduced previously.
@@ 6.3 Quarkonium production at electron-positron colliders
To complete the picture of quarkonium production across different collider types, we present in this section predictions for cross sections in $e^{+}e^{-}$ collisions. Measurements from the Belle and BaBar $B$ factories provide one of the most precise data on quarkonium production at $e^{+}e^{-}$ colliders, including both inclusive processes such as $e^{+}e^{-}\to J/\psi+X$ BaBar:2001lfi ; Belle:2001lqi ; Belle:2009bxr and $e^{+}e^{-}\to J/\psi+c\bar{c}+X$ Belle:2002tfa ; Belle:2009bxr , and exclusive processes such as $e^{+}e^{-}\to J/\psi+\eta_{c}$ Belle:2002tfa ; Belle:2004abn ; BaBar:2005nic ; Belle:2023gln and $e^{+}e^{-}\to J/\psi+\gamma$ BaBar:2003waz . In the following, we present examples of both inclusive and exclusive quarkonium production processes in sections 6.3.1 and 6.3.2, respectively.
@@ 7.1 Positronium production
Nonetheless, to demonstrate the theoretical capabilities of our framework, we present in this section several cross-section predictions for positronium production in $e^{+}e^{-}$ and $pp$ collisions. Here, we restrict ourselves to ground-state positronium with the principal number $N=1$, although MG5_aMC also supports calculations for excited leptonia with $N>1$. Depending on the relative spin orientation of the electron and positron, two distinct ground-state configurations can form. Para-positronium ($\mathit{Ps}_{0}$) with total angular momentum $J=0$ (corresponding to the Fock state $n={}^{1}S_{0}$) arises from antialigned spins, where the spin vectors point in opposite directions. Conversely, if the spins are aligned parallel, ortho-positronium ($\mathit{Ps}_{1}$) with total angular momentum $J=1$ (corresponding to the Fock state $n={}^{3}S_{1}$) is formed.
@@
Table 13: Cross-section predictions for para- ($\mathit{Ps}_{0}$) and ortho-positronium ($\mathit{Ps}_{1}$) production at $\sqrt{s}=13$ TeV in $pp$ collisions and $\sqrt{s}=10.58\,\mathrm{GeV}$ in $e^{+}e^{-}$ collisions, based on the setup described in section 4. Additional fiducial cuts of $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ ($p_{T,\gamma}>0.5\,\mathrm{GeV}$ and $|\eta_{\gamma}|<5$) are applied to final-state jets (photons) in the computations for ortho-positronium production. All cross sections are rescaled according to the $\alpha(0)$ scheme. The numbers in parentheses indicate the numerical uncertainties from the MC phase-space integration.
@@
where we impose fiducial jet cuts of $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ in addition to the common setup to ensure infrared safety. We obtain a cross section of $10.489(2)\,\mathrm{fb}$, which might make experimental detection conceivable. The corresponding $p_{T,\mathit{Ps}_{1}}$ distribution is shown in figure 5(a), where the uncertainty band represents a 7-point scale variation with $\mu_{R/F}\in\{\mu_{0}/2,\mu_{0},2\mu_{0}\}$ around the central scale $\mu_{0}=H_{T}/2$, obeying $1/2\leq\mu_{R}/\mu_{F}\leq 2$.
@@
Before turning to other leptonia, we analyse the numerical stability and accuracy of the predictions obtained with MG5_aMC by comparing them to analytic computations in the decoupling limit of the $Z$ and Higgs bosons. In this limit, the para-positronium cross section for the process in Eq. (93) reduces to the expression
@@
Figure 6: Associated positronium-production cross sections for (a) para-positronium plus photon and (b) ortho-positronium plus photon in $e^{+}e^{-}$ collisions. The upper panels show the total cross sections as functions of the $e^{+}e^{-}$ c.m. energy, obtained with MG5_aMC (blue dots) and from analytic expressions (orange line). The lower panels display the ratio of the two results, normalised to the analytic predictions. Error bars on the MG5_aMC points indicate statistical MC uncertainties.
Figure 6 compares the numerical predictions from MG5_aMC with these analytic results as a function of the ratio of the $e^{+}e^{-}$ c.m. energy to the positronium mass. To obtain the infinitely heavy $Z$ and $H$ boson limit in MG5_aMC, we use the prompts:

### v2

@@ 1 Introduction
The main motivation of this work is to develop and ultimately provide the community with an all-encompassing tool for bound-state production studies in NRQCD and non-relativistic QED (NRQED) Caswell:1985ui . The goal is to deliver reliable, efficient, and fast automated state-of-the-art computations, with NLO accuracy based on an extended Frixione–Kunszt–Signer (FKS) infrared-divergence subtraction formalism Frixione:1995ms ; Frixione:1997np ; AH:2024ueu planned for near-term implementation. This work lays the foundation for that endeavour and falls within the aforementioned category of process-independent parton-level event generators, designed to simulate processes involving one or more quarkonia and/or leptonia. As mentioned, our implementation is embedded within the well-established MG5_aMC framework, which is widely used in the high-energy physics community, thereby extending its capabilities to include S-wave quarkonium and leptonium production. To the best of our knowledge, no other publicly available tool currently provides similar functionality for leptonium processes.
@@ 2 Theoretical framework
The short-distance cross section ${\rm d}\hat{\sigma}({\cal I}_{1}{\cal I}_{2}\rightarrow(C_{1}C_{2})[n]+X_{p})$ can be obtained from the production amplitude of $C_{1}$ and $C_{2}$ by projecting it onto the specific quantum state $n$, following section 2 of ref. AH:2024ueu . As mentioned previously, the quantum state $n$ is specified by its $S$, $L$, $J$ and $C$ quantum numbers. For leptonium ($\mathcal{B}=\mathcal{L}$), only the colour singlet configuration ($C=1$) is possible.
@@ 5 Benchmark processes and validation
In this section, we benchmark our LO implementation by comparing its numerical results for various observables across multiple processes in $e^{+}e^{-}$ and $pp$ collisions with the corresponding results from HELAC-Onia Shao:2012iz ; Shao:2015vga . The comparisons include processes with single and multiple quarkonium final states, as well as quarkonium production in association with SM elementary particles. Our aim is to validate our calculations and assess their consistency with other established tools. This benchmarking provides an important cross check, ensuring the reliability of our results for the applications discussed in the following two sections.
@@ 6 Quarkonium production
To highlight the broad applicability of our implementation across different collision systems, we present cross sections for quarkonium production in proton-proton ($pp$) collisions (section 6.1), electron-proton ($e^{-}p$) collisions (section 6.2), and electron-positron ($e^{+}e^{-}$) collisions (section 6.3).
@@
We emphasise that the broad range of processes considered in what follows is intended to demonstrate the versatility of our implementation; the numerical values presented should not be directly compared with experimental measurements, where available. As previously stated, our current implementation is restricted to S-wave production at LO and does not, for instance, account for feeddown effects where they occur. Nevertheless, we provide a sufficiently detailed exposition of the results to allow the reader to reproduce all numerical values shown using MG5_aMC with the same parameter inputs. In addition, we make our analyses scripts available as auxiliary files in the supplemental material accompanying the arXiv submission.
@@ 6.1.4 $J/\psi$ production in association with an open charm-anticharm pair
The LHCb experiment has observed the associated production of $J/\psi+D$ LHCb:2012aiv and $\Upsilon+D$ LHCb:2015wvu , with results indicating that SPS contributions alone are insufficient to describe the data, thus requiring the inclusion of DPS effects. Motivated by this strong phenomenological relevance, numerous theoretical studies have been dedicated to these processes Berezhnoy:2015jga ; Likhoded:2015fdr ; Karpishkov:2019vyt ; Shao:2020kgj ; Chernyshev:2023qea .
@@
Figure 1: Transverse momentum ($p_{T,\mathcal{Q}}$) distributions for $J/\psi$ production in association with $c\bar{c}$ in $pp$ collisions at $\sqrt{s}=13$ TeV. Panels (b-d) correspond to different $J/\psi$ Fock states, while panel (a) shows the total contribution obtained by combining them. The blue curve represents the fixed-order LO (fLO) computation from MG5_aMC, and the orange curve shows the MG5_aMC LO events combined with PS from Pythia v8.315 (LO+PS). Both include their respective 7-point scale-variation uncertainty bands. The lower panels display the ratio to the fLO contribution in each case.
@@
In figure 1, we compare the $J/\psi$ transverse momentum spectra from the fixed-order LO (fLO) calculation and the LO events matched to PS (LO+PS). The $p_{T,J/\psi}$ spectrum including all three Fock states is shown in figure 1(a), while figures 1(b)1(c), and 1(d) display the spectra for the individual $J/\psi\!\left[{}^{3}S_{1}^{[1]}\right$], $J/\psi\!\left[{}^{1}S_{0}^{[8]}\right$], and $J/\psi\!\left[{}^{3}S_{1}^{[8]}\right$] Fock states, respectively. We observe that PS effects broaden the $p_{T,J/\psi}$ distributions in the $J/\psi\!\left[{}^{3}S_{1}^{[1]}\right$] and $J/\psi\!\left[{}^{1}S_{0}^{[8]}\right$] channels, while they narrow the transverse-momentum distribution in the $J/\psi\!\left[{}^{3}S_{1}^{[8]}\right$] channel. These opposite shifts from the two spin-triplet Fock states partially compensate each other in the total (figure 1(a)), resulting in a moderate overall distortion of the $p_{T,J/\psi}$ spectrum. Unlike the processes $pp\to J/\psi\!\left[{}^{3}S_{1}^{[8]},\,^{1}S_{0}^{[8]},\,^{3}P_{J}^{[8]}\right]+j+X$, the cross sections for these processes at small $p_{T,J/\psi}$ are finite. In the former case, one would need to impose cuts, and the effect of showering in presence of such cuts would be more significant Lansberg:2024zap .
@@ 6.1.5 Inclusive quarkonium-pair production
In contrast, $J/\psi+\Upsilon$ production in hadron-hadron collisions has now been measured at both the Tevatron by D0 D0:2015dyx and at the LHC by LHCb LHCb:2023qgu . The main motivation for studying charmonium plus bottomonium production is to probe colour-octet quarkonium production mechanisms Ko:2010xy ; Lansberg:2015lva ; Lansberg:2020rft , DPS Likhoded:2015zna ; Shao:2016wor , and potential new exotic hadrons such as fully-heavy tetraquarks ($b\bar{b}c\bar{c}$). Unlike $J/\psi$-pair or $\Upsilon$-pair production, the SPS colour-singlet process $pp\to J/\psi\!\left[{}^{3}S_{1}^{[1]}\right]+\Upsilon\!\left[{}^{3}S_{1}^{[1]}\right]+X$ vanishes in QCD at LO ($\mathcal{O}(\alpha_{s}^{4}$)) due to quantum-number conservation. In fact, the lowest-order contribution in QCD is loop-induced at $\mathcal{O}(\alpha_{s}^{6})$ Shao:2016wor . This suppression of the SPS colour-singlet channel makes DPS contributions more prominent and increases the relative importance of colour-octet mechanisms in the SPS channel.
@@
Table 9 shows our results for the production of charmonium or bottomonium pairs, while table 10 presents the production of mixed charmonium–bottomonium pairs as well as $B_{c}^{\pm}$ pairs. All cross sections are computed at $\mathcal{O}(\alpha_{s}^{4})$. For charmonium or bottomonium pairs, the cross sections are dominated by colour-singlet configurations when both quarkonia are of the same type, e.g., in the production of two $\eta_{c}$ ($\eta_{b}$) or two $J/\psi$ ($\Upsilon$) mesons. By contrast, at LO in QCD, processes involving one pseudoscalar and one vector quarkonium of the same flavour, such as $\eta_{c}+J/\psi$ ($\eta_{b}+\Upsilon$), are typically dominated by configurations involving at least one colour-octet state. This is because, due to C parity, the pure colour-singlet partonic channels $gg\to(Q\bar{Q})\!\left[{}^{1}S_{0}^{[1]}\right]+(Q\bar{Q})\!\left[{}^{3}S_{1}^{[1]}\right$] and $q\bar{q}\to(Q\bar{Q})\!\left[{}^{1}S_{0}^{[1]}\right]+(Q\bar{Q})\!\left[{}^{3}S_{1}^{[1]}\right$] are forbidden in QCD. This also explains why the cross sections for pseudoscalar-vector final states are smaller than those of the other two configurations in table 9.
@@
We again emphasise the broad physics potential of quarkonium-pair production studies for probing the interplay between SPS and DPS interactions. However, the numerical values presented in this work account only for the SPS contribution at LO, using generic cuts and LDME values, for illustration purposes.
@@ 6.1.6 Inclusive triple $J/\psi$ production
To showcase the capabilities of our implementation, we consider triple $J/\psi$ production in $pp$ collisions. This serves as an illustrative example of a non-trivial three-quarkonium final state, proposed in refs. Shao:2019qob ; dEnterria:2016ids ; dEnterria:2017yhd as a probe of triple parton scattering (TPS). Recently measured by CMS CMS:2021qsn , this marks the first experimental measurement of TPS, providing new insight into multi-parton scattering dynamics. The LO ($\mathcal{O}(\alpha_{s}^{6})$) total cross section in QCD at $\sqrt{s}=13\,\mathrm{TeV}$ is
@@ 6.2 Quarkonium production in electron-proton collisions
At electron-proton colliders, inelastic quarkonium production can occur via either DIS or photoproduction processes. In the following two subsections, we consider both types of collisions and compute cross sections for charmonium production. We retain the same SM input parameters as discussed in section 4, and here specify the relevant collider settings, which were not introduced previously.
@@ 6.3 Quarkonium production at electron-positron colliders
To complete the picture of quarkonium production across different collider types, we present in this section numbers for cross sections in $e^{+}e^{-}$ collisions. Measurements from the Belle and BaBar $B$ factories provide one of the most precise data on quarkonium production at $e^{+}e^{-}$ colliders, including both inclusive processes such as $e^{+}e^{-}\to J/\psi+X$ BaBar:2001lfi ; Belle:2001lqi ; Belle:2009bxr and $e^{+}e^{-}\to J/\psi+c\bar{c}+X$ Belle:2002tfa ; Belle:2009bxr , and exclusive processes such as $e^{+}e^{-}\to J/\psi+\eta_{c}$ Belle:2002tfa ; Belle:2004abn ; BaBar:2005nic ; Belle:2023gln and $e^{+}e^{-}\to J/\psi+\gamma$ BaBar:2003waz . In the following, we present examples of both inclusive and exclusive quarkonium production processes in sections 6.3.1 and 6.3.2, respectively.
@@ 7.1 Positronium production
Nonetheless, to demonstrate the theoretical capabilities of our framework, we present in this section several cross sections for positronium production in $e^{+}e^{-}$ and $pp$ collisions. Here, we restrict ourselves to ground-state positronium with the principal number $N=1$, although MG5_aMC also supports calculations for excited leptonia with $N>1$. Depending on the relative spin orientation of the electron and positron, two distinct ground-state configurations can form. Para-positronium ($\mathit{Ps}_{0}$) with total angular momentum $J=0$ (corresponding to the Fock state $n={}^{1}S_{0}$) arises from antialigned spins, where the spin vectors point in opposite directions. Conversely, if the spins are aligned parallel, ortho-positronium ($\mathit{Ps}_{1}$) with total angular momentum $J=1$ (corresponding to the Fock state $n={}^{3}S_{1}$) is formed.
@@
Table 13: Cross-section results for para- ($\mathit{Ps}_{0}$) and ortho-positronium ($\mathit{Ps}_{1}$) production at $\sqrt{s}=13$ TeV in $pp$ collisions and $\sqrt{s}=10.58\,\mathrm{GeV}$ in $e^{+}e^{-}$ collisions, based on the setup described in section 4. Additional fiducial cuts of $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ ($p_{T,\gamma}>0.5\,\mathrm{GeV}$ and $|\eta_{\gamma}|<5$) are applied to final-state jets (photons) in the computations for ortho-positronium production. All cross sections are rescaled according to the $\alpha(0)$ scheme. The numbers in parentheses indicate the numerical uncertainties from the MC phase-space integration.
@@
where we impose fiducial jet cuts of $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ in addition to the common setup to ensure infrared safety. We obtain a cross section of $9.483(2)\,\mathrm{fb}$, which might make experimental detection conceivable. The corresponding $p_{T,\mathit{Ps}_{1}}$ distribution is shown in figure 5(a), where the uncertainty band represents a 7-point scale variation with $\mu_{R/F}\in\{\mu_{0}/2,\mu_{0},2\mu_{0}\}$ around the central scale $\mu_{0}=H_{T}/2$, obeying $1/2\leq\mu_{R}/\mu_{F}\leq 2$.
@@
Before turning to other leptonia, we analyse the numerical stability and accuracy of the values obtained with MG5_aMC by comparing them to analytic computations in the decoupling limit of the $Z$ and Higgs bosons. In this limit, the para-positronium cross section for the process in Eq. (93) reduces to the expression
@@
Figure 6: Associated positronium-production cross sections for (a) para-positronium plus photon and (b) ortho-positronium plus photon in $e^{+}e^{-}$ collisions. The upper panels show the total cross sections as functions of the $e^{+}e^{-}$ c.m. energy, obtained with MG5_aMC (blue dots) and from analytic expressions (orange line). The lower panels display the ratio of the two results, normalised to the analytic reference values. Error bars on the MG5_aMC points indicate statistical MC uncertainties.
Figure 6 compares the numerical outputs from MG5_aMC with these analytic results as a function of the ratio of the $e^{+}e^{-}$ c.m. energy to the positronium mass. To obtain the infinitely heavy $Z$ and $H$ boson limit in MG5_aMC, we use the prompts: